# Concentrated Monte Carlo sampling for local observables in quantum spin chains

### v1

@@ I Introduction
Monte Carlo(MC) simulations have become a standard tool for estimating observables in statistical physics by sampling configurations from a target measure. Its reach spans statistical mechanics [binder1997applications, newman1999monte, landau2021guide], molecular electronic-structure theory [ceperley1980ground, foulkes2001quantum, austin2012quantum], quantum lattice systems [hirsch1985two, sandvik1991quantum, sandvik2010computational, capogrosso2008monte, sandvik2007evidence, varney2009quantum, mazurenko2017cold], and neural quantum states [carleo2017solving, carleo2019machine, hibat2020recurrent, sharir2020deep, zhang2025paths, gravina2025neural, van2025many, sinibaldi2024timedependentneuralgalerkinmethod, WangPoletti2025]; and has also led to the development of specialized MC algorithms that improve sampling efficiency and accuracy for classical and quantum many-body systems, such as Markov chain Monte Carlo (MCMC) [metropolis1953equation, hastings1970monte], stochastic series expansion algorithm [sandvik1999stochastic, syljuaasen2002quantum], worm algorithm [prokof2001worm, prokof1998worm], auxiliary-field quantum Monte Carlo [blankenbecler1981monte, zhang1997constrained, zhang2003quantum], path-integral Monte Carlo [feynman2018statistical, ceperley1995path, boninsegni2006worm], variational Monte Carlo [foulkes2001quantum, becca2017quantum, toulouse2016introduction]. With MC approaches, one can estimate large sums or integrals by sample averages over configurations drawn from a target distribution, and the standard error decreases as $1/\sqrt{N_{s}}$, where $N_{s}$ is the number of samples, independent of the system’s dimension.
@@ III Ground State of Tilted Ising Model
Since the purpose is to evaluate the effectiveness of the sampling, we evaluate very accurately the ground state and its obervables using the matrix product state (MPS) method [white1992density, schollwock2011density] to produce both a reliable wave-function to sample from, and the ground truth for value for the observables. We will focus on the single-site observable $\langle\sigma^{z}_{i}\rangle$, which represents the expectation value of local magnetization in the $z$-direction, as well as two-site observables $\langle\sigma^{z}_{i}\sigma^{z}_{i+1}\rangle$, corresponding to the spin-spin correlation in the $z$-direction. For each quantity, we perform $R=10$ independent runs and compute the mean absolute error $\epsilon(\langle O_{loc}\rangle)=\frac{1}{R}\sum_{r}|\langle O_{loc}\rangle-\langle O_{loc}\rangle_{\text{ex}}|$, where $\langle O_{loc}\rangle_{\text{ex}}$ is the exact value, here from MPS. Specifically, we focus on a 20-spin TIM with $2^{20}$ configurations, which provides a sufficiently large Hilbert space such that the sampling remains nontrivial.
Figure 2: Performance of CMCS and MCMC for the different ground states of 1D 20 spins tilted Ising model, presented for different local region sizes $\ell$. The panels show the mean absolute error $\epsilon$ of $\langle\sigma^{z}_{i}\rangle$ as a function of the site index $i$, where the exact values are obtained by MPS. The system is in (a) the paramagnetic phase with $h_{x}=10J,\;h_{z}=J/2$, (b) the critical region with $h_{x}=0.95J,\;h_{z}=J/2$, (c) the antiferromagnetic phase with $h_{x}=J/2,h_{z}=J/2$ and (d) the ferromagnetic phase with $J<0,h_{x}=|J|/2,\;h_{z}=|J|/2$. $\ast$ represents CMCS $\ell=2$ with sites 10-11, $N_{l}=4$ and $100\leq{N_{u}}\leq 2000$; $\lozenge$ represents CMCS $\ell=3$ with sites 9-11, $N_{l}=8$ and $192\leq{N_{u}}\leq 4000$; $\bigtriangleup$ represents CMCS $\ell=4$ with sites 9-12, $N_{l}=16$ and $352\leq{N_{u}}\leq 8000$.
@@
In Fig. 2, we benchmark the performance of CMCS against conventional MCMC across different ground states by plotting the mean absolute error of $\langle\sigma^{z}_{i}\rangle$ as a function of the site index $i$. In all panels, the local region used in CMCS is placed in the center of the chain and its size $\ell$ is varied as indicated. We can see that, except near criticality, CMCS outperforms MCMC within the chosen central local region, whereby local observables evaluated inside the local region exhibit significantly smaller errors. Specifically, in panel (a), we consider the system to be in the paramagnetic phase, where the probability distribution in the $\sigma^{z}$ basis is close to uniform and the true value of $\langle\sigma^{z}_{i}\rangle$ is very small but not exactly 0. In this case, MCMC suffers from statistical noise of order $1/\sqrt{N_{s}}$ that can overwhelm the signal, so the small but finite expectation value cannot be statistically resolved from zero at the current sample size. In contrast, the same near-uniform distribution means that most environment configurations have very similar weights, and from the perspective of the CMCS algorithm, the local observable is therefore weakly sensitive to the specific environment configurations. This increases the effective sample size and suppresses the noise, so that these near-zero values of local observable are accurately resolved by CMCS. In the ordered phases shown in panels (c) and (d), the probability weight is concentrated around a few ordered configurations. For local-update MCMC this leads to long autocorrelation times and the Markov chain is easily trapped near local minima. Obtaining reliable results with MCMC in this case would therefore require much longer chains and many more samples, although it can be improved with importance sampling techniques [alain2016variancereductionsgddistributed, katharopoulos2019samplescreatedequaldeep, misery2025lookingelsewhereimprovingvariational, yang2020deep]. CMCS is well suited to such highly concentrated distributions, as it still explores the configurations within a local region, yielding precise local observables even when the initial number of Monte Carlo samples is relatively small. Around the critical region shown in panel (b), the correlation length becomes large and local observables become sensitive to distant parts of the system. Insufficient sampling of the environment configurations can then have a strong impact on the local observables. Under these conditions, CMCS generally does not outperform conventional MCMC.
@@ IV Thermal States
Where $\theta$ is an angular variables, and $\pmb{S}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ is the vector of spin-1 operators at the site $i$ [parkinson1988s, binder2020low]. The model can be tuned by varying the angle $\theta$: when $\pi/4<\theta<\pi/4$, the system is in the Haldane phase; when $\pi/4<\theta<\pi/2$, the system is in the critical phase; when $\pi/2\leq\theta<5\pi/4$, the system is in the ferromagnetic phase, and when $\theta=\pi/4,\pi/2$, the system is in the phase transition point. When $\theta=0$, the Haldane characterizes the antiferromagnetic-Heisenberg (AFH) state, and when $\theta=\arctan(1/3)$, the system is described by a topological order, which is the Affleck-KennedyLieb-Tasaki (AKLT) state [affleck1987rigorous, affleck1988valence]. The AKLT state is a valence-bond solid on a spin-1 chain. At each site, the spin-1 is decomposed into two spin-1/2s, which are projected onto the symmetric ($S=1$) triplet subspace; on every nearest-neighbor bond, one spin-1/2 from each site forms a singlet. The AKLT state is an exact MPS with bond dimension $D=2$, in which connected two-point correlators decay exponentially with distance. For open chains, the unpaired edge spin-1/2s give rise to characteristic edge states.
@@ VI Acknowledgements
We acknowledge fruitful discussions with E. Khatami, E. Merali and R.T. Scalettar. We also acknowledge the support of the Ministry of Education, Singapore, under the grant MOE-T2EP50123-0017, and from the Centre for Quantum Technologies grant CQT_\_SUTD_\_2025_\_01. Part of the numerical work was performed at the National Supercomputing Centre, Singapore [nscc].

### v2

@@ I Introduction
Monte Carlo(MC) simulations have become a standard tool for estimating observables in statistical physics by sampling configurations from a target measure. Its reach spans statistical mechanics [7, 27, 23], molecular electronic-structure theory [14, 17, 5], quantum lattice systems [21, 31, 34, 11, 33, 41, 24], and neural quantum states [13, 12, 20, 36, 47, 18, 40, 37, 42]; and has also led to the development of specialized MC algorithms that improve sampling efficiency and accuracy for classical and quantum many-body systems, such as Markov chain Monte Carlo (MCMC) [25, 19], stochastic series expansion algorithm [32, 38], worm algorithm [29, 30], auxiliary-field quantum Monte Carlo [9, 45, 46], path-integral Monte Carlo [16, 15, 10], variational Monte Carlo [17, 6, 39]. With MC approaches, one can estimate large sums or integrals by sample averages over configurations drawn from a target distribution, and the standard error decreases as $1/\sqrt{N_{s}}$, where $N_{s}$ is the number of samples, independent of the system’s dimension.
@@ III Ground State of Tilted Ising Model
Since the purpose is to evaluate the effectiveness of the sampling, we evaluate very accurately the ground state and its obervables using the matrix product state (MPS) method [43, 35] to produce both a reliable wave-function to sample from, and the ground truth for value for the observables. We will focus on the single-site observable $\langle\sigma^{z}_{i}\rangle$, which represents the expectation value of local magnetization in the $z$-direction, as well as two-site observables $\langle\sigma^{z}_{i}\sigma^{z}_{i+1}\rangle$, corresponding to the spin-spin correlation in the $z$-direction. For each quantity, we perform $R=10$ independent runs and compute the mean absolute error $\epsilon(\langle O_{loc}\rangle)=\frac{1}{R}\sum_{r}|\langle O_{loc}\rangle-\langle O_{loc}\rangle_{\text{ex}}|$, where $\langle O_{loc}\rangle_{\text{ex}}$ is the exact value, here from MPS. Specifically, we focus on a 20-spin TIM with $2^{20}$ configurations, which provides a sufficiently large Hilbert space such that the sampling remains nontrivial.
Figure 2: Performance of CMCS and MCMC for the different ground states of 1D 20 spins tilted Ising model, presented for different local region sizes $\ell$. The panels show the mean absolute error $\epsilon$ of $\langle\sigma^{z}_{i}\rangle$ as a function of the site index $i$, where the exact values are obtained by MPS. The system is in (a) the paramagnetic phase with $h_{x}=10J,\;h_{z}=J/2$, (b) the critical region with $h_{x}=0.95J,\;h_{z}=J/2$, (c) the antiferromagnetic phase with $h_{x}=J/2,h_{z}=J/2$ and (d) the ferromagnetic phase with $J<0,h_{x}=|J|/2,\;h_{z}=|J|/2$. $\circ$ represents MC sampling with $N_{s}=10^{4}$; $\ast$ represents CMCS $\ell=2$ with sites 10-11, $N_{l}=4$ and $100\leq{N_{u}}\leq 2000$; $\lozenge$ represents CMCS $\ell=3$ with sites 9-11, $N_{l}=8$ and $192\leq{N_{u}}\leq 4000$; $\bigtriangleup$ represents CMCS $\ell=4$ with sites 9-12, $N_{l}=16$ and $352\leq{N_{u}}\leq 8000$.
@@
In Fig. 2, we benchmark the performance of CMCS against conventional MCMC across different ground states by plotting the mean absolute error of $\langle\sigma^{z}_{i}\rangle$ as a function of the site index $i$. In all panels, the local region used in CMCS is placed in the center of the chain and its size $\ell$ is varied as indicated. We can see that, except near criticality, CMCS outperforms MCMC within the chosen central local region, whereby local observables evaluated inside the local region exhibit significantly smaller errors. Specifically, in panel (a), we consider the system to be in the paramagnetic phase, where the probability distribution in the $\sigma^{z}$ basis is close to uniform and the true value of $\langle\sigma^{z}_{i}\rangle$ is very small but not exactly 0. In this case, MCMC suffers from statistical noise of order $1/\sqrt{N_{s}}$ that can overwhelm the signal, so the small but finite expectation value cannot be statistically resolved from zero at the current sample size. In contrast, the same near-uniform distribution means that most environment configurations have very similar weights, and from the perspective of the CMCS algorithm, the local observable is therefore weakly sensitive to the specific environment configurations. This increases the effective sample size and suppresses the noise, so that these near-zero values of local observable are accurately resolved by CMCS. In the ordered phases shown in panels (c) and (d), the probability weight is concentrated around a few ordered configurations. For local-update MCMC this leads to long autocorrelation times and the Markov chain is easily trapped near local minima. Obtaining reliable results with MCMC in this case would therefore require much longer chains and many more samples, although it can be improved with importance sampling techniques [4, 22, 26, 44]. CMCS is well suited to such highly concentrated distributions, as it still explores the configurations within a local region, yielding precise local observables even when the initial number of Monte Carlo samples is relatively small. Around the critical region shown in panel (b), the correlation length becomes large and local observables become sensitive to distant parts of the system. Insufficient sampling of the environment configurations can then have a strong impact on the local observables. Under these conditions, CMCS generally does not outperform conventional MCMC. Similar results are obtained when measuring $\langle\sigma^{x}_{i}\rangle$, see App. A.
@@ IV Thermal States
Where $\theta$ is an angular variables, and $\pmb{S}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ is the vector of spin-1 operators at the site $i$ [28, 8]. The model can be tuned by varying the angle $\theta$: when $\pi/4<\theta<\pi/4$, the system is in the Haldane phase; when $\pi/4<\theta<\pi/2$, the system is in the critical phase; when $\pi/2\leq\theta<5\pi/4$, the system is in the ferromagnetic phase, and when $\theta=\pi/4,\pi/2$, the system is in the phase transition point. When $\theta=0$, the Haldane characterizes the antiferromagnetic-Heisenberg (AFH) state, and when $\theta=\arctan(1/3)$, the system is described by a topological order, which is the Affleck-KennedyLieb-Tasaki (AKLT) state [3, 2]. The AKLT state is a valence-bond solid on a spin-1 chain. At each site, the spin-1 is decomposed into two spin-1/2s, which are projected onto the symmetric ($S=1$) triplet subspace; on every nearest-neighbor bond, one spin-1/2 from each site forms a singlet. The AKLT state is an exact MPS with bond dimension $D=2$, in which connected two-point correlators decay exponentially with distance. For open chains, the unpaired edge spin-1/2s give rise to characteristic edge states.
@@ VI Acknowledgements
We acknowledge fruitful discussions with E. Khatami, E. Merali and R.T. Scalettar. We also acknowledge the support of the Ministry of Education, Singapore, under the grant MOE-T2EP50123-0017, and from the Centre for Quantum Technologies grant CQT_\_SUTD_\_2025_\_01. Part of the numerical work was performed at the National Supercomputing Centre, Singapore [1].
Appendix
Appendix A Estimation of generic operators
Here, we estimate the expectation value of a generic operator $O$ with respect to wave function $|\psi\rangle=\sum_{\pmb{x}}\psi(\pmb{x})\,|\pmb{x}\rangle$ expressed in the computational basis,. It is given by
$\langle O\rangle=$ $\frac{\langle\psi|O|\psi\rangle}{\langle\psi|\psi\rangle}$ (11)
$=$ $\sum_{\pmb{x}}\frac{\psi^{*}(\pmb{x})}{\langle\psi|\psi\rangle}\sum_{\pmb{x}^{\prime}}O_{\pmb{x},\pmb{x}^{\prime}}\psi(\pmb{x}^{\prime})$
$=$ $\sum_{\pmb{x}}\frac{|\psi^{*}(\pmb{x})|^{2}}{\langle\psi|\psi\rangle}\sum_{\pmb{x}^{\prime}}O_{\pmb{x},\pmb{x}^{\prime}}\frac{\psi(\pmb{x^{\prime}})}{\psi(\pmb{x})}$
$=$ $\sum_{\pmb{x}}P(\pmb{x})O_{loc}(\pmb{x})$
In the CMCS scheme, we renormalize the wave function on the set $\mathcal{U}$ of unique configurations generated by the chosen local region. It reads
$\widetilde{\psi}(\pmb{u})=\frac{\psi(\pmb{u})}{\sqrt{\sum_{\pmb{u}}\psi^{*}(\pmb{u})\psi(\pmb{u})}}$ (12)
and the renormalized probability is
$\widetilde{P}(\pmb{u})=\frac{|\psi(\pmb{u})|^{2}}{\sum_{\pmb{u}}\psi^{*}(\pmb{u})\psi(\pmb{u})}$ (13)
Thus, the expectation value of the operator in CMCS method is
$\langle O\rangle=$ $\sum_{\pmb{u}\in\mathcal{U}}\widetilde{P}(\pmb{u})\sum_{\pmb{u}^{\prime}}O_{\pmb{u},\pmb{u}^{\prime}}\frac{\psi(\pmb{u^{\prime}})}{\psi(\pmb{u})}$ (14)
For diagonal operators ($O_{\pmb{x,x^{\prime}}}=O(\pmb{x})\delta_{\pmb{x,x^{\prime}}}$), only the configuration $\pmb{x^{\prime}}=\pmb{x}$ contributes, and no extra configurations need to be included; off-diagonal operators, instead, require that all connected states $\pmb{x}^{\prime}$ are contained in $\mathcal{U}$. Fig. 6 shows the mean absolute error of $\langle\sigma_{i}^{x}\rangle$ as an example off-diagonal observable. Since site $i$ must lie in the local region, the local region is shifted along the chain as $i$ is varied. We observe that its performance is comparable to that of the diagonal observable $\langle\sigma_{i}^{z}\rangle$, depicted in Fig. 2.
Figure 6: Comparison of the mean absolute error of operator $\langle\sigma^{x}_{i}\rangle$ between MCMC and CMCS for the ground state of 1D 20 spins tilted Ising model using different local-region sizes $\ell$. The system is in (a) the paramagnetic phase with $h_{x}=10J,h_{z}=J/2$, (b) the critical region with $h_{x}=0.95J,h_{z}=J/2$, (c) the antiferromagnetic phase with $h_{x}=J/2,h_{z}=J/2$, and (d) the ferromagnetic phase with $J<0,h_{x}=|J|/2,h_{z}=|J|/2$. The local region is shifted along the chain as the site $i$ is varied. For each local region, the measurement site where $\langle\sigma_{i}^{x}\rangle$ is measured is chosen as follows: for $\ell=2$, we use the first site of the region; for $\ell=3$, we use the central site of the region; and for $\ell=4$, we use the two central sites of the region (measurements at the boundary sites of the local region are also considered). $\circ$ represents MC sampling with $N_{s}=10^{4}$; $\ast$ represents CMCS $\ell=2$ ; $\lozenge$ represents CMCS $\ell=3$ ; $\bigtriangleup$ represents CMCS $\ell=4$.